A copper wire of cross sectional area 0.01 mm2 is used to prepare a resistance of 1kohm The resistivity of copper is 1.7x10x8the length of the wire is
1
Expert's answer
2019-07-19T14:10:08-0400

The resistance of a wire can be calculated as follows:


"R = \rho \frac{l}{S},"

where \rho is the resistivity of a wire, l is its length and S is a cross-sectional area.

Hence,


"l = \frac{R S}{\rho}"

Substituting the numerical values, we obtain:


"l=\frac{10^3 \Omega \cdot 10^{-2} \cdot 10^{-6} m^2}{1.7 \cdot 10^{-8} \Omega \cdot m} \approx 588 \, m"

Answer: 588 m.